OpenStudy (anonymous):

What is the slope, and what is the y-intercept of the line? y=5/4x+9 slope=[ ]/[ ] y-intercept=[ ]

OpenStudy (mayankdevnani):

\[\large \bf y=mx+c\] where, \[\large \bf m=slope \[\large \bf c=y-intercept\]

OpenStudy (mayankdevnani):

\[\large \bf m=slope ~and~c=y-intercept\]

OpenStudy (mayankdevnani):

So,compare with your question, \[\large \bf y=\frac{5}{4}x+9\] \[\large \bf y=mx+c\]

OpenStudy (jojo4eva):

so the slope would be 5/4 and the y intercept would be 9 based off of the info i have given you beforehand

OpenStudy (mayankdevnani):

can you indentify m(slope) and c(y-intercept) ?? @Samus_Aran

OpenStudy (anonymous):

slope 5 ans y-intercept 9?

OpenStudy (mayankdevnani):

wrong answer for slope but y-intercept is correct

OpenStudy (mayankdevnani):

figure out again ? BETTER LUCK NEXT TIME :)

OpenStudy (anonymous):

x for slope?

OpenStudy (anonymous):

oh wait 5/4

OpenStudy (mayankdevnani):

correct ^^

OpenStudy (mayankdevnani):

good job :)

OpenStudy (anonymous):

thanks @mayankdevnani

OpenStudy (mayankdevnani):

welcome :)